OpenStudy (anonymous):

in a fastfood chain, four cheeseburger and two large drinks cost 180. 2 large drinks cost 5 more than a cheeseburger. what is the cost of the cheeseburger? the drink?

OpenStudy (anonymous):

C= # of cheesburgers D=# of drinks 4C+2D=180 2D=C+5 --> rearrange: C=2D-5 plug one equation into the other: 4(2D-5)+2D=180 8D-20+2D=180 10D-20=180 add 20 to both sides: 10D=200 divide both sides by 10 D=20 plug back into the original equation: C=2D-5 C=2(20)-5 C=40-5 C=35 The cheeseburger costs 35 while the drink costs 20.

OpenStudy (anonymous):

thank you